Skip to Search

 
diagonal spacer

Information Technology Staff Meeting Notes
March 20, 2009

Present: Thomas Bielicke, Cynthia Bowker, Laurie Brennan, Will Chaney, Logan Clapp, Matthew Daley, Debra Fetzer, Jason Fleming, Winston Harris, Michael Jay, Paul McDonough, Bobbie Meng, Tom Minton, Chris Nicolich, Amy Polk, Mark Sullivan, Evan Wack, Stephen Williams
Announcements
  1. Kanapaha festival is this weekend.
  2. Basketball bracket discussion.
  3. Dean Russell to attend our staff meeting on March 27.
  4. Employee Recognition Committee will fund an event in recognition of the IT Dept's extra effort to quell the public machine virus outbreak several months back. We have decided to have a catered lunch during one of our Friday staff meetings, since we will all be together at that time anyways.
Ongoing IT Updates
Equipment Received/Installed/Repaired
  1. Will and Logan are putting together an order for the possible expansion of the SAN. DLC may expand their service to the entire state. We're putting together a scenario to handle it. DLC is already offering the PALMM service. It would be modeled on the DLOC toolkit. DLC would be looking to add .5-2 FTE, yet the project could save money in the long run.
  2. We are also putting together a list of replacement workstations for staff and public.
Support Issues
  1. UFAD migration - the bulk of the workstations are done. Only machines that are not in use or are used for test purposes on Smatherslib have not migrated.
Training
  1. Cyber Self Defense - March 23.
Hardware
  1. Laptop encryption - uninstalling VPN had no effect - will pass those tickets on to Will.
  2. Admin rights - adding to a separate OU - Matthew Mariner's machine. The OU for DLC and Ariel was moved to coordinate with the building.
  3. Acrobat vulnerability - must download the latest version, 9.0.1, to fix. Will require hands-on in the public areas.
  4. Education-Journalism VLAN changes - not complete. Still need to do the laptops in the Education classroom as well as three other machines.
  5. Xerox is making changes today to accommodate the VLAN changes. CNS (not campus CNS) was over at Education, hammering and drilling.
  6. Auto play - if it is disabled, there will be no pop-up. Will put it on hold for now.
  7. SP3 problems - Amy thinks it's fixed. If you find any machines with the symptoms, let her know.
  8. SCCM moving to UFAD - will have to wait for SQL2K8 to have it up and running. May take up to a month. We're working on it.
  9. EPO - still can't see the public machines. They are updating, but reporting to the wrong server.
  10. Green Team request to power off monitors after 10 minutes of inactivity - the policy is ready to go. We're waiting for the AGRSSR group to say it's acceptable and then we'll need to notify staff. Start date: April 20.
Hardware Support Unit
  1. Inventory - finished, with the exception of the IT Dept. We just wrapped up in Special Collections - there's a lot of quirkiness there. Need to follow up on policy override. Need to check and see if the policy is applying in DLC as well. Discussion on rebuilding all of Special Collections public machines.
  2. NT2R backups - offline static image - no need to back up. All dailies are backed up to NT16, except FS1. Keep doing that way until we find a better solution.
  3. IT Dept inventory scheduled for Monday. Will need assistance in the server room. Also, need key to the server rack in MSL.
Software Projects Status
  1. ARL stats - Read Only to be done at the end of the month.
  2. FilmLog is done - it can be removed from the list.
  3. ITS load - done.
  4. Root cert - prices may vary, depending on the company. Need to evaluate. Bill advises to go with a major company, to assure mobile devices will be recognized. Ok, we'll go with Thawte.
  5. Chris has been cleaning up NT4. There are still a few applications linked to it. Should be done by next Friday.
  6. SQL2K8 upgrade - license to be ordered. Server is configured for it. Need best practice for the software installation.
  7. New test server for apps is ready to go.
  8. ILLIAD billing taskforce to meet next week. Stephen is compiling statistics.
  9. ILLIAD - HSC & Law - has to be coordinated with OCLC. Moving from one symbol to multi symbol may be a difference in workstation configuration - they use templates, which may need to be modified based on the version of Office. Michelle wants Office 2K7 to be installed for ILL. It needs to be in writing - she should submit a ticket.
  10. Request from HR to collect degree information for faculty and staff on the Green Monster. They want staff to enter it for themselves - crossover into Gator Scholar. Skills & Training database - should contain the data.
  11. Stephen Williams is putting together a quick java script to detect mobile devices for the Web to deliver a better version.
  12. Request for weekly Patron Update for FCLA to include barcode info - from Jim Stevens. Chris is looking onto it.
  13. Tom is working on OCR toolkit automation.
  14. Mark - working on Google map searching results in DLC. It displays polygons and is a challenge to format on the page.
  15. UFDC - new version - more browsing capabilities - Mark has been working on it.
  16. Bridges has produced the new patron affiliation, but it can't be seen. It's not functioning as advertised. Winston will ask to have it exposed. Warren should be available to help.
ALEPH
  1. They were supposed to upgrade to Linux VMWare, but it was held up by CNS with cabling. Unknown how much it will hold up the upgrade to version 19. Upgrade is scheduled to be available in test by April and up by July 1.
  2. Last Endeca test load for DLC today - includes thumbnails.
Web Projects
  1. Tom has been getting lots of requests to promote stuff on the home page - different collections, sub-guides and databases. May have one or two spots that would rotate promotions.
  2. Development - log-in box for Friends of the Library. Tom could make a special log-in box, but on the Proxy server it could be confusing. Tom will talk to Barbara Hood.
  3. Today is Matthew's last day until May. He will transfer his Web tickets to Tom. Remember not to make him a Secondary. Self-Check and Aleph Print problems to be handled by Evan.
DLC
  1. Drives installed last Friday, mounted arrays and added space. JPEG2K not helped - it has USB drive issues.
  2. Logan is still working on the SAN - mounting space - it is not 100% yet.
  3. Greenstone migration is contingent upon the purchase of new space.
  4. Greenstone test VM to be created by Mark. Once we have space, we can move Greenstone. One of the proposals for space is just for DLC. The other is for DLC and PALMM.
  5. Plans to reorganize DLC SAN space - on hold pending space purchase.
Students and Department
  1. IT Dept staff relocating - Tom and Matthew will move to room 419 Library East. It will also house a usability lab - no longer to be used as an IT Training room. Michelle Crump will be moving to the room next door. All of the equipment currently stored in room 419 will have to be moved down to the IT Dept.
IT Security Management
  1. Meeting held.

 

University of Florida Home Page